5. Your broker promises that if you give her $15,000 today she will return $30,000 to you in five years. To the nearest percent, what annual interest rate is being offered?

6. How much money would you have to put away at the end of each year to have $1,250,000 when you retire 42 years from now if you can earn 5% on your money?

7. How much can be accumulated if $500 per month is deposited for the next 35 years, beginning next month, if the account earns 6% APR (i.e. 0.5% per month).

8. What is the price of a $1,000 par value bond with a 6% annual coupon that matures in 7 years if the interest rate is 8%?

9. What is the yield to maturity of a 6-year, 6% bond that currently sells for $1,049.17?

10. How much would an investor expect to pay for a $1,000 par value bond with a 6% annual coupon that matures in 7 years if the interest rate is 3%?
Answer #1

15% 6 $9,243.39 7 $712,355.15 8 $895.87 9 5.03% 10 $1,186.91 5 =RATE(5,,-15000,30000) 6 =-PMT(5%, 42.1250000) 7 =FV(6%/12,35*